OpenStudy (anonymous):

Given that; a=(2,-3), b=(3,4), c=(-1,5), find the number. (a-b)*(b+c)

OpenStudy (anonymous):

just so you know this section is with the dot product.

OpenStudy (anonymous):

\[a - b = <2,3> - <3,4> = <2-3, 3-4> = <-1,-1>\] \[b + c = <3,4> + <-1,5> = <3-1, 4+5> = <2,9>\] \[(a-b)\cdot(b+c) = <-1,-1>\cdot<2,9>\]

OpenStudy (anonymous):

\[=-1\cdot 2 + -1\cdot 9 = -2 -9 = -11\]

OpenStudy (anonymous):

(-1,-7)*(2,9) =>-65 is the answer.

OpenStudy (anonymous):

polpak u missed in the first step i guess...

OpenStudy (anonymous):

Oh, woops.. I missed the -3.

OpenStudy (anonymous):

In any case the process is the same.

OpenStudy (anonymous):

I have 3 instead of -3.

OpenStudy (anonymous):

Thanks you guys!!